OpenStudy (anonymous):

solve ⎷‾13x+13‾ = x+1 (13x+13 is under a factor sign)

OpenStudy (turingtest):

what, you mean a radical sign? \[\sqrt{13x+13}=x+1\]

OpenStudy (anonymous):

yes thats it!

OpenStudy (anonymous):

you want to solve for x?

OpenStudy (anonymous):

yes. its X= either 1 or 2 answers

OpenStudy (turingtest):

sure, square both sides to get rid of the sign 13x+13=(x+1)^2=(x+1)(x+1)=x^2+2x+1 combine like terms and set equal to zero: x^2-11x-12=0 (x-12)(x+1)=0 x={-1,12}

OpenStudy (turingtest):

follow all that?

OpenStudy (anonymous):

GOT IT! thx so so much :)

OpenStudy (anonymous):

\[\left( \sqrt{13x+13} \right)^{2} = \left( x+1 \right)^{2}\] \[x ^{2}+2x+1=13x+13\] \[x ^{2}-11x-12\] x=(-1,12)

OpenStudy (anonymous):

champion is right